During a chase with the Joker, Batman drove 55 km due East then cut 55° to the right and drove another 30 km. What is Batman’s displacement?


1
Expert's answer
2021-02-28T07:36:00-0500

"\vec{s}=(55,0)+(30\cos{55},30\sin{55})\\\vec{s}=(72.2,24.6)\ km\\s=\sqrt{72.2^2+24.6^2}=76.3\ km\\\theta=\arctan{\frac{24.6}{72.2}}=18.8\degree\text{ S of E}"